# FreeBird.jl: An Extensible Toolbox for Simulating Interfacial Phase Equilibria

### v1

@@ 3.1.3 LJ Surfaces
Finally, we examine the (111) surface of a face-centered cubic LJ solid, building on previous NS benchmarks. 29, 30 We model a five-layer slab containing 80 fixed particles (16 per layer) and four mobile adsorbates, corresponding to a maximum coverage of . The simulation box has periodic boundaries in the in-plane directions ( Å, Å) and a non-periodic upper boundary at Å where simple reflection is applied if a particle moves beyond this limit. Sampling parameters are as follows: NS ran for 50,000 steps with 320 walkers; MC cooled from 2000 to 100 K in 100 K increments; and WL sampled energies from -50.0 to -57.06 eV, slightly above the ground-state energy (-57.07 eV) (Figure 8).
This LJ(111) quarter-coverage (q.c.) system is expected to display a broad peak at (surface condensation) and a sharper peak at (two-dimensional adsorbate ordering). 29 Figure 8 shows that all three methods yield consistent transition temperatures in units of : 0.833 (NS), 0.862 (MC), and 0.849 (WL) for the condensation transition; and 0.265 (NS), 0.259 (MC), and 0.247 (WL) for the ordering transition.
@@ 3.1.4 Discussion of Atomistic Demonstrations
For the simple atomistic systems examined, NS, MC, and WL produced consistent configurational statistics and thermodynamic properties. The agreement among the three methods, and with published benchmarks, 62, 63, 29 confirms that they are correctly implemented within FreeBird.jl. This unified framework allows verification of sampling outputs by cross-checking across methods using a consistent call structure, making it straightforward for users to apply multiple algorithms to their systems of interest.
@@
Because all sampling methods use the same random-walk and energy-evaluation routines, the total number of energy evaluations provides a consistent probe of computational cost (Table 1). Note that these costs vary with both system complexity and algorithm choice. Table 1 is not intended as a direct performance comparison, since the methods differ in their intrinsic overheads (e.g., number of walkers for NS, temperature grids for MC, and number of bins plus flatness criterion for WL), and each can be further tuned to balance accuracy and cost. The reported values serve only as references for results obtained using straightforward, representative parameter settings for each algorithm.
@@
Although direct comparisons are challenging due to differing objectives and strategies, the sampling methods exhibit distinct trade-offs. NS is well-suited for generating configurations that span a wide energy range, from gas-like states to ground-state structures, making it effective for identifying candidate transition regions. Its stepwise nature limits parallelism compared with approaches such as parallel tempering, 43 but its walker-removal rate can be tuned to emphasize either exploratory breadth or resolution. 42 This flexibility makes NS a strong first-pass strategy for locating regions of interest for subsequent refinement with other algorithms. MC is advantageous when properties at specific temperatures are of primary interest. WL, though computationally demanding for continuous systems, 64 can yield low-variance estimates once convergence is reached, due to its comprehensive sampling of configurations finely spaced in energy. In practice, the methods can complement each other: NS can identify temperature ranges for MC and guide energy bounds for WL, while MC energy histograms can inform WL binning resolution within those bounds. As a modular toolbox, FreeBird.jl supports not only method selection but also coordination across algorithms. The three demonstrations underscore the framework’s flexibility, accommodating varying numbers of components, interaction schemes, and physical constraints. As the codebase evolves, it will continue to support increasingly complex and realistic systems.
@@ 3.2.2 Three-Dimensional Lattice Models
We next examine a three-dimensional primitive cubic lattice with four occupied sites. The bottom layer represents an adsorbent surface, while the upper two layers act as a fluid phase into which particles may desorb, emulating surface condensation. The same lattice Hamiltonian is used as in Section 3.2.1, except that eV is assigned to occupied sites in the fluid layers to isolate the effect of fluid-phase introduction. Only nn and nnn interactions are considered. Simulation parameters match those of the two-dimensional model. Exact enumeration of the full configuration space (194,580 configurations) again reveals an order–disorder transition at lower temperatures (, nearly identical to the transition in the 2D case), favoring the square ground state. In contrast to the two-dimensional system, however, the three-layer model also shows a broad peak at higher temperatures (), corresponding to a surface condensation transition. As shown in Figure 9, all three methods capture both features with closely matching transition temperatures: 0.319 (NS), 0.259 (MC), and 0.327 (WL) for the order–disorder transition; and 1.629 (NS), 1.637 (MC), and 1.637 (WL) for the surface condensation transition.
@@ 3.2.3 Discussion of Lattice Demonstrations
NS poses additional challenges. In lattice models, discrete energy degeneracies produce plateaus in the NS likelihood–energy profile. 65, 66, 67 To break these degeneracies, we follow the approach of Murray et al. and recent BraWl implementations, adding to each configuration’s energy a uniformly distributed perturbation much smaller than the typical level spacing (e.g., eV). 65, 68 With single-site random-walk moves, lattice NS shows sharply reduced acceptance rates as the energy ceiling approaches the ground-state energy. Introducing cluster or collective moves, or increasing the live-set size, can restore sampling efficiency. 69 Even so, WL remains the most efficient method for these systems.
@@ 4 Conclusions and Outlook
Julia, as a high-level, just-in-time compiled language, significantly lowers the barrier to developing complete, high-performance computational chemistry packages. 70 The availability of high-quality standard libraries and community packages in the Julia ecosystem, particularly those from JuliaMolSim, has been instrumental in the rapid development of FreeBird.jl. Examples include AtomsBase.jl 71 for handling atomic and cell properties, AtomsIO.jl 72 and ExtXYZ.jl 73 for reading and writing atomic structures, and the Julia standard libraries Threads and Distributed for straightforward parallelization, which substantially improves FreeBird.jl’s performance and enables execution on HPC systems. Contributors to this open-source package span a range of career stages, from undergraduate students to established computational scientists, underscoring its accessibility to both users and developers. Finally, FreeBird.jl exemplifies the adoption of modern software engineering practices to enhance productivity, sustainability, and reproducibility, serving as a template for the development of contemporary scientific software.
Future developments of FreeBird.jl can be categorized into algorithmic and implementation enhancements. Algorithmically, integrating the existing sampling methods with ensembles beyond the canonical, such as the semi-grand-canonical ensemble (to allow composition changes) and the grand-canonical ensemble (for chemical potential control), will expand the range of systems and properties relevant to materials discovery. Additionally, while the current move sets are adequate for many systems, more sophisticated proposals (e.g., cluster or collective moves) are often necessary near critical points or in systems with slow dynamics to ensure efficient sampling. 69, 74 From an implementation perspective, enabling FreeBird.jl for efficient GPU computing is essential for scaling to exascale platforms. This capability can be combined with machine-learning interatomic potentials for energy evaluations, which naturally exploit GPU acceleration. Overall, FreeBird.jl functions not only as a performant tool for sampling computations, but also as an integrative interface linking diverse components of molecular simulation, and as a versatile platform for advancing computational chemistry and materials science.

### v3

@@ 3.1.3 LJ Surfaces
Finally, we examine the (111) surface of a face-centered cubic LJ solid, building on previous NS benchmarks. 32, 33 We model a five-layer slab containing 80 fixed particles (16 per layer) and four mobile adsorbates, corresponding to a maximum coverage of . The simulation box has periodic boundaries in the in-plane directions ( Å, Å) and a non-periodic upper boundary at Å where simple reflection is applied if a particle moves beyond this limit. Sampling parameters are as follows: NS ran for 50,000 steps with 320 walkers; MC cooled from 2000 to 100 K in 100 K increments; and WL sampled energies from -50.0 to -57.06 eV, slightly above the ground-state energy (-57.07 eV) (Figure 8).
This LJ(111) quarter-coverage (q.c.) system is expected to display a broad peak at (surface condensation) and a sharper peak at (two-dimensional adsorbate ordering). 32 Figure 8 shows that all three methods yield consistent transition temperatures in units of : 0.833 (NS), 0.862 (MC), and 0.849 (WL) for the condensation transition; and 0.265 (NS), 0.259 (MC), and 0.247 (WL) for the ordering transition. These findings are in exact agreement with our earlier study employing an independent nested sampling implementation. 32
@@ 3.1.4 Discussion of Atomistic Demonstrations
For the simple atomistic systems examined, NS, MC, and WL produced consistent configurational statistics and thermodynamic properties. The agreement among the three methods, and with published benchmarks, 73, 76, 32, 75 confirms that they are correctly implemented within FreeBird.jl. This unified framework allows verification of sampling outputs by cross-checking across methods using a consistent call structure, making it straightforward for users to apply multiple algorithms to their systems of interest.
@@
Because all sampling methods use the same random-walk and energy-evaluation routines, the total number of energy evaluations provides a consistent probe of computational cost (Table 1). The reported values range from a few million to several billion evaluations, illustrating the order-of-magnitude differences in computational effort between algorithms. Note that these costs vary with both system complexity and algorithm choice. Table 1 is not intended as a direct performance comparison, since the methods differ in their intrinsic overheads (e.g., number of walkers for NS, temperature grids for MC, and number of bins plus flatness criterion for WL), and each can be further tuned to balance accuracy and cost. The reported values serve only as references for results obtained using straightforward, representative parameter settings for each algorithm.
@@
Although direct comparisons are challenging due to differing objectives and strategies, the sampling methods exhibit distinct trade-offs. NS is well-suited for generating configurations that span a wide energy range, from gas-like states to ground-state structures, making it effective for identifying candidate transition regions. Its stepwise nature limits parallelism compared with approaches such as parallel tempering, 46 but its walker-removal rate can be tuned to emphasize either exploratory breadth or resolution. 45 This flexibility makes NS a strong first-pass strategy for locating regions of interest for subsequent refinement with other algorithms. MC is advantageous when properties at specific temperatures are of primary interest. WL, though computationally demanding for continuous systems, 78 can yield low-variance estimates once convergence is reached, due to its comprehensive sampling of configurations finely spaced in energy. In practice, the methods can complement each other: NS can identify temperature ranges for MC and guide energy bounds for WL, while MC energy histograms can inform WL binning resolution within those bounds. As a modular toolbox, FreeBird.jl supports not only method selection but also coordination across algorithms. The three demonstrations underscore the framework’s flexibility, accommodating varying numbers of components, interaction schemes, and physical constraints. As the codebase evolves, it will continue to support increasingly complex and realistic systems.
@@ 3.2.2 Three-Dimensional Lattice Models
We next examine a three-dimensional primitive cubic lattice with four occupied sites. The bottom layer represents an adsorbent surface, while the upper two layers act as a fluid phase into which particles may desorb, emulating surface condensation. The same lattice Hamiltonian is used as in Section 3.2.1, except that eV is assigned to occupied sites in the fluid layers to isolate the effect of fluid-phase introduction. Only nn and nnn interactions are considered. Simulation parameters match those of the two-dimensional model. Exact enumeration of the full configuration space (194,580 configurations) again reveals an order–disorder transition at lower temperatures (, nearly identical to the transition in the 2D case), favoring the square ground state. In contrast to the two-dimensional system, however, the three-layer model also shows a broad peak at higher temperatures (), corresponding to a surface condensation transition. As shown in Figure 9, all three methods capture both features with closely matching transition temperatures: 0.319 (NS), 0.259 (MC), and 0.327 (WL) for the order–disorder transition; and 1.629 (NS), 1.637 (MC), and 1.637 (WL) for the surface condensation transition. Compared to the 2D case, the computational effort for all methods was increased to ensure convergence. The number of NS iterations was raised from 10,000 to 30,000; MC equilibration and sampling steps at each temperature were increased from 25,000 to 100,000; and WL retained the same number of energy bins (since the potential-energy landscape is unchanged), but the number of MC proposals per attempt to reduce was increased from 100 to 1,000 to account for the larger number of microstates. These settings were used to generate the results shown in Figure 9(b).
@@ 3.2.3 Discussion of Lattice Demonstrations
NS poses additional challenges. In lattice models, discrete energy degeneracies produce plateaus in the NS likelihood–energy profile. 79, 80, 81 To break these degeneracies, we follow the approach of Murray et al. and recent BraWl implementations, adding to each configuration’s energy a uniformly distributed perturbation much smaller than the typical level spacing (e.g., eV). 79, 82 With single-site random-walk moves, lattice NS shows sharply reduced acceptance rates as the energy ceiling approaches the ground-state energy. As shown in Figure 9, 10,000 NS samples are required for the 2D lattice to establish the correct distribution for constructing an accurate partition function, exceeding the total number of microstates in the system. Introducing cluster or collective moves, or increasing the live-set size, can restore sampling efficiency. 83 Even so, WL remains the most efficient method for these systems.
@@ 4 Conclusions and Outlook
Julia, as a high-level, just-in-time compiled language, significantly lowers the barrier to developing complete, high-performance computational chemistry packages. 84 The availability of high-quality standard libraries and community packages in the Julia ecosystem, particularly those from JuliaMolSim, has been instrumental in the rapid development of FreeBird.jl. Examples include AtomsBase.jl 85 for handling atomic and cell properties, AtomsIO.jl 86 and ExtXYZ.jl 87 for reading and writing atomic structures, and the Julia standard libraries Threads and Distributed for straightforward parallelization, which substantially improves FreeBird.jl’s performance and enables execution on HPC systems. Contributors to this open-source package span a range of career stages, from undergraduate students to established computational scientists, underscoring its accessibility to both users and developers. Finally, FreeBird.jl exemplifies the adoption of modern software engineering practices to enhance productivity, sustainability, and reproducibility, serving as a template for the development of contemporary scientific software.
Future developments of FreeBird.jl can be categorized into algorithmic and implementation enhancements. Algorithmically, integrating the existing sampling methods with ensembles beyond the canonical, such as the semi-grand-canonical ensemble (to allow composition changes) and the grand-canonical ensemble (for chemical potential control), will expand the range of systems and properties relevant to materials discovery. Additionally, while the current move sets are adequate for many systems, more sophisticated proposals (e.g., cluster or collective moves) are often necessary near critical points or in systems with slow dynamics to ensure efficient sampling. 83, 88 From an implementation perspective, enabling FreeBird.jl for efficient GPU computing is essential for scaling to exascale platforms. This capability can be combined with machine-learning interatomic potentials for energy evaluations, which naturally exploit GPU acceleration. Overall, FreeBird.jl functions not only as a performant tool for sampling computations, but also as an integrative interface linking diverse components of molecular simulation, and as a versatile platform for advancing computational chemistry and materials science.